# Equivariance and partial observations in Koopman operator theory for partial differential equations

### v2

@@ Abstract
The Koopman operator has become an essential tool for data-driven analysis, prediction and control of complex systems. The main reason is the enormous potential of identifying linear function space representations of nonlinear dynamics from measurements. This equally applies to ordinary, stochastic, and partial differential equations (PDEs). Until now, with a few exceptions only, the PDE case is mostly treated rather superficially, and the specific structure of the underlying dynamics is largely ignored. In this paper, we show that symmetries in the system dynamics can be carried over to the Koopman operator, which allows us to massively increase the model efficacy. Moreover, the situation where we only have access to partial observations—i.e., measurements, as is very common for experimental data—has not been treated to its full extent, either. Moreover, we address the highly-relevant case where we cannot measure the full state, such that alternative approaches such as delay coordinates have to be considered. We derive rigorous statements on the required number of observables in this situation, based on embedding theory. We present numerical evidence using various numerical examples including the wave equation and the Kuramoto-Sivashinsky equation.
@@ 1 Introduction
Due to these reasons, there has been an increasing interest in the scientific community to develop and improve methods to infer and predict dynamical systems from data. Popular examples are the Sparse Identification of Nonlinear Dynamics [BPK16], statistical approaches such as the Mori-Zwanzig framework [CHK00], or techniques based on deep learning [VBW+18], also using physical knowledge [RPK19]. Another approach that has been particularly successful in the past decade is the Koopman operator framework [Koo31, RMB+09, Mez13]. The driving force behind this renaissance is twofold: (1) The Koopman operator yields a linear representation of nonlinear dynamical systems, thus giving us access to powerful techniques from linear systems theory, and (2) the advances in numerical approximation (here in the form of the Extended Dynamic Mode Decomposition (EDMD) [WKR15, KKS16, KNP+20]) now allow us to identify the Koopman operator from measurements using simple linear regression. As a consequence, the Koopman operator has been studied extensively for data-driven analysis [BBKK22], coarse graining [BNC18, KNP+20, NKS21, NKBC21] and control [PBK15, KM18a, PK19, POR20, SWP+23, BGSW23] of a very large number of applications including neuroscience, social systems, robotics, and fluid dynamics. Recent results on finite-data error bounds can be found in [ZZ23, NPP+23, PSW+24a, BBL+23, KPS+24].
@@ 2 Koopman operator for PDEs
Here, $y(x,t)$ is the space- and time-dependent system state, $t\in\mathbb{R}_{\geq 0}$ is the time and $x\in\Omega$ the spatial coordinate, where $\Omega\subset\mathbb{R}^{n}$ is the spatial domain with boundary $\partial\Omega$. For simplicity of notation, we will restrict our analysis to one-dimensional spatial domains. However, extensions to higher dimensions are possible in a straight-forward manner. Moreover, the state at time $t$ is an element of a function space with appropriate differentiability (e.g., a Sobolev space $y(\cdot,t)\in H^{p}(\Omega)$) and $\mathcal{N}:D(\mathcal{N})\rightarrow H^{p}(\Omega)$ (with $D(\mathcal{N})\subset H^{p}(\Omega)$ being the domain of $\mathcal{N}$) is a nonlinear partial differential operator describing the dynamics of the system. We will for now assume simple rectangular domains and periodic boundary conditions (BCs). The system’s flow map $\Phi^{\tau}:L^{2}(\Omega)\to L^{2}(\Omega)$ is defined by
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(\mathcal{K}f\right)(y)=\lim_{\tau\to 0}\frac{\left(\mathcal{K}^{\tau}f\right)(y)-f(y)}{\tau}\qquad\forall y\in\mathcal{A}.}$ (4)
@@
The goal is then to approximate the compression $P_{\mathbb{V}}\mathcal{K}^{\tau}|_{\mathbb{V}}$, where $\mathbb{V}=\operatorname{span}\{\psi_{j}|j\in\{1,\ldots,\ell\}\}$, and $P_{\mathbb{V}}$ is the projection onto $\mathbb{V}$. Using time series data $[\Psi(y_{0}),\Psi(y_{1}),\ldots,\Psi(y_{m})]$, where $y_{i}=y(\cdot,i\tau)$ and $\Psi=[\psi_{1}^{\top},\cdots,\psi_{\ell}^{\top}]^{\top}$, we can now simply use linear regression to find the best fit matrix $K^{\tau}$:
@@
A very similar regression problem can be formulated to approximate the generator $\mathcal{K}$ via $K$ [KNP+20]. In both cases, it can be shown that this matrix converges to the Galerkin projection of $\mathcal{K}^{\tau}$ in the infinite data limit $m\rightarrow\infty$ [WKR15, KKS16], and to the true Koopman operator when additionally $\ell\rightarrow\infty$ [KM18b]. Moreover, finite-data error bounds can be found in [Mez22, ZZ23, NPP+23, PSW+24a, KPS+24], using either i.i.d. or ergodic sampling. Most of them have until now considered ordinary or stochastic differential equations. A more general setting can be found in [PSW+24b], which covers the PDE case as well.
@@ 3.3 Equivariant Koopman operators
In many situations, $\theta$ is a Gaussian kernel—i.e., $\theta=\exp(-\|s-x\|/\alpha)$, with kernel bandwidth $\alpha\in\mathbb{R}^{>0}$—that somewhat “localizes” $y_{t}$ around the center $s$ (even though not in a strict sense, of course). Now, fixing $s$, we can define the observable function
@@
Consider a PDE of the general form (1) with periodic boundary conditions, where $\mathcal{N}$ does not explicitly depend on space $x$ or time $t$ and is thus equivariant under translations in $x$ (and $t$) in view of the periodic boundary conditions. Then, the Koopman operator associated with (1) and the observable $f_{s}$ as defined in (9) is also equivariant under the same group action.
@@ 4.2 Relation between the Core Dynamical System and the underlying PDE
Let $\mathcal{A}\subset\mathcal{Y}$ be a compact and invariant set with upper box counting dimension $\dim(\mathcal{A})=d$ and thickness exponent $\sigma$. Choose an integer $q>2(1+\sigma)d$, and suppose further that the set $\mathcal{A}_{p}$ of $p$-periodic points of $\Phi^{\tau}$ satisfies $\dim(\mathcal{A}_{p})<p/(2+2\sigma)$ for $p=1,...,q$. Then for almost every (in the sense of prevalence) Lipschitz map $g:\mathcal{Y}\rightarrow\mathbb{R}$ the delay observation map $f:=D_{q}[g,\Phi^{\tau}]:\mathcal{Y}\rightarrow\mathbb{R}^{q}$ defined by
$y\mapsto\begin{bmatrix}g(y)&g(\Phi^{\tau}(y))&\ldots&g(\Phi^{(q-1)\tau}(y)\end{bmatrix}^{\top}$
is one-to-one on $\mathcal{A}$. The same holds for a set of $q$ distinct observables $g_{1},\ldots,g_{q}:\mathcal{Y}\to\mathbb{R}$, i.e.,
$f=\begin{bmatrix}g_{1}(y)&\ldots&g_{q}(y)\end{bmatrix}^{\top}.$
@@
The central message of the above theorem is that we can draw a close connection between the observable $f$ in the Koopman setting (Definition 1 and Remark 2) and the observation map $g$ in the embedding framework. The key statement for our purposes is that if the system dynamics of $\Phi^{\tau}$ is restricted to a compact set $\mathcal{A}$ with a finite dimension $d$, then we need to have at least $q>2d$ distinct measurements $g$—which jointly form the Koopman observable $f$—to obtain a one-to-one correspondence between $\Phi^{\tau}$ and $\varphi^{\tau}$:
$\Phi^{\tau}=E\circ\varphi^{\tau}\circ D_{q}[g,\Phi^{\tau}]=E\circ\varphi^{\tau}\circ f,$
@@ 5 Combining equivariance and partial measurements
In fact, if we were able to obtain an exact finite-dimensional approximation of $\hat{\mathcal{K}}^{\tau}$, then we would quickly run into inconsistencies, as the same matrix would be a predictor for $(z_{1},z_{2},z_{3})$ and for $(z_{2},z_{3},z_{4})$. A quick calculation then shows that the prediction of any $z_{i}$ cannot depend on any other $z_{j}$, which would mean that $K^{\tau}$ is just a diagonal matrix. Our interpretation of this dilemma is that inconsistencies cannot be avoided, as in the generic situation, any finite-dimensional approximation is inexact, meaning that the local Koopman models have to be globally inconsistent or yield trivial dynamics. Note that this situation is different for more sophisticated observable functions such as Fourier modes or Koopman eigenfunctions—these, however, do not allow for shift-equivariance as we consider here.
@@
As the above approach only works for control inputs that enter the original system in a purely linear fashion [NPP+23], an alternative approach is to accept that a purely linear approach may be too much to ask for. Instead, a coupling can be achieved by predicting the next state, project from $\Psi(z)$ onto the coordinates $z$, and then lift again for each system. Due to the project-then-lift step, we “synchronize” the local systems, at the cost of obtaining nonlinear dynamics. Similar observations in terms of trading the linearity for accuracy have been made in, e.g., [CALG23]. Note that if we use kernel EDMD with a radial basis function kernel, then this approach possesses a relation to Gaussian Process models.

### v3

@@ Abstract
The Koopman operator has become an essential tool for data-driven analysis, prediction and control of complex systems. The main reason is the enormous potential of identifying linear function space representations of nonlinear dynamics from measurements. This equally applies to ordinary, stochastic, and partial differential equations (PDEs). Until now, with a few exceptions only, the PDE case is mostly treated rather superficially, and the specific structure of the underlying dynamics is largely ignored. In this paper, we show that symmetries in the system dynamics can be carried over to the Koopman operator, which allows us to significantly increase the model efficacy. Moreover, the situation where we only have access to partial observations—i.e., measurements, as is very common for experimental data—has not been treated to its full extent, either. Moreover, we address the highly-relevant case where we cannot measure the full state, where alternative approaches (e.g., delay coordinates) have to be considered. We derive rigorous statements on the required number of observables in this situation, based on embedding theory. We present numerical evidence using various numerical examples including the wave equation and the Kuramoto-Sivashinsky equation.
@@ 1 Introduction
Due to these reasons, there has been an increasing interest in the scientific community to develop and improve methods to infer and predict dynamical systems from data. Popular examples are the Sparse Identification of Nonlinear Dynamics [BPK16], statistical approaches such as the Mori-Zwanzig framework [CHK00], or techniques based on deep learning [VBW+18], also using physical knowledge [RPK19]. Another approach that has been particularly successful in the past decade is the Koopman operator framework [Koo31, RMB+09, Mez13]. The driving force behind this renaissance is twofold: (1) The Koopman operator yields a linear representation of nonlinear dynamical systems, thus giving us access to powerful techniques from linear systems theory, and (2) the advances in numerical approximation (here in the form of the Extended Dynamic Mode Decomposition (EDMD) [WKR15, KKS16, KNP+20]) now allow us to identify the Koopman operator from measurements using simple linear regression. As a consequence, the Koopman operator has been studied extensively for data-driven analysis [BBKK22], coarse graining [BNC18, KNP+20, NKS21, NKBC21] and control [PBK15, KM18a, PK19, POR20, SWP+23, BGSW24] of a very large number of applications including neuroscience, social systems, robotics, and fluid dynamics. Recent results on finite-data error bounds can be found in [ZZ23, NPP+23, PSW+24, BBL+23, KPS+24].
@@ 2 Koopman operator for PDEs
Here, $y(x,t)$ is the space- and time-dependent system state, $t\in\mathbb{R}_{\geq 0}$ is the time and $x\in\Omega$ the spatial coordinate, where $\Omega\subset\mathbb{R}^{n}$ is the spatial domain with boundary $\partial\Omega$. For simplicity of notation, we will restrict our analysis to one-dimensional spatial domains. However, extensions to higher dimensions are possible in a straight-forward manner. Moreover, the state at time $t$ is an element of a function space with appropriate differentiability (e.g., a Sobolev space $y(\cdot,t)\in H^{p}(\Omega)$) and $\mathcal{N}:D(\mathcal{N})\rightarrow H^{p}(\Omega)$ (with $D(\mathcal{N})\subset H^{p}(\Omega)$ being the domain of $\mathcal{N}$) is a nonlinear partial differential operator describing the dynamics of the system. We will for now assume periodic boundary conditions (BCs). The system’s flow map $\Phi^{\tau}:L^{2}(\Omega)\to L^{2}(\Omega)$ is defined by
@@
$\left(\mathcal{K}f\right)(y)=\lim_{\tau\to 0}\frac{\left(\mathcal{K}^{\tau}f\right)(y)-f(y)}{\tau}\qquad\forall y\in\mathcal{A}.$ (4)
@@
The goal is then to approximate the compression $P_{\mathbb{V}}\mathcal{K}^{\tau}|_{\mathbb{V}}$, where $\mathbb{V}=\operatorname{span}\{\psi_{j}|j\in\{1,\ldots,\ell\}\}$, and $P_{\mathbb{V}}$ is the projection onto $\mathbb{V}$. Using time series data $\{\Psi(y_{k})\}_{k=0}^{m}$, where $y_{i}=y(\cdot,i\tau)$ and $\Psi=[\psi_{1}^{\top},\cdots,\psi_{\ell}^{\top}]^{\top}$, we can now simply use linear regression to find the best fit matrix $K^{\tau}$:
@@
A very similar regression problem can be formulated to approximate the generator $\mathcal{K}$ via $K$ [KNP+20]. In both cases, it can be shown that this matrix converges to the Galerkin projection of $\mathcal{K}^{\tau}$ in the infinite data limit $m\rightarrow\infty$ [WKR15, KKS16], and to the true Koopman operator when additionally $\ell\rightarrow\infty$ [KM18b]. Moreover, finite-data error bounds can be found in [Mez22, ZZ23, NPP+23, PSW+24, KPS+24], using either i.i.d. or ergodic sampling. Most of them have until now considered ordinary or stochastic differential equations. A more general setting can be found in [PSB+24], which covers the PDE case as well.
@@ 3.3 Equivariant Koopman operators
In many situations, $\theta$ is a Gaussian kernel—i.e., $\theta=\exp(-\|s-x\|^{2}/\alpha)$, with kernel bandwidth $\alpha\in\mathbb{R}^{>0}$—that somewhat “localizes” $y_{t}$ around the center $s$ (even though not in a strict sense, of course). Now, fixing $s$, we can define the observable function
@@
Consider a PDE of the general form (1) with periodic boundary conditions, where $\mathcal{N}$ does not explicitly depend on space $x$ or time $t$ and is thus equivariant under translations in $x$ (and $t$) in view of the periodic boundary conditions. Further assume that the observable $f_{s}$ (of the form as defined in (9)) is periodic as well. Then, the Koopman operator associated with (1) is also equivariant under the same group action.
@@
Note that the specific form of the observable (9) excludes global transformations such as the Fourier transform. For these, the symmetry transformation has more sophisticated implications that are less intuitive, which we emphasize in the following example. For a more thorough discussion of the requirements for symmetry-preserving observables, see [HPN+24].
Example 3.2 (Fourier observable).
Let us assume that—instead of Eq. (9)—our observable $f$ is defined as the (complex) Fourier transform for some fixed frequency $\omega$, i.e.,
$f(y)=\mathcal{F}\{y\}(\omega)=\int_{\mathbb{R}^{n}}y(x)e^{-i\omega x}\,\mathrm{d}x,$
and that our group action $g$ introduces a translation: ${g}\left[{y}\right](x)=y(x-g)$. We then find
$f({g}\left[{y}\right])$ =$=\mathcal{F}\{{g}\left[{y}\right]\}(\omega)=\int_{\mathbb{R}^{n}}{g}\left[{y}\right](x)e^{-i\omega x}\,\mathrm{d}x=\int_{\mathbb{R}^{n}}y(x-g)e^{-i\omega x}\,\mathrm{d}x$
$=\int_{\mathbb{R}^{n}}y(x)e^{-i\omega(x+g)}\,\mathrm{d}x=e^{-i\omega g}\int_{\mathbb{R}^{n}}y(x)e^{-i\omega x}\,\mathrm{d}x$
$=e^{-i\omega g}f(y)={g}\left[{f}\right](y).$
We thus see that the group action applied to $f$ results in the multiplication with a complex number.
@@ 4.2 Relation between the Core Dynamical System and the underlying PDE
Let $\mathcal{A}\subset\mathcal{Y}$ be a compact and invariant set with upper box counting dimension $\dim(\mathcal{A})=d$ and thickness exponent $\sigma$. Choose an integer $q>2(1+\sigma)d$, and suppose further that the set $\mathcal{A}_{p}$ of $p$-periodic points of $\Phi^{\tau}$ satisfies $\dim(\mathcal{A}_{p})<p/(2+2\sigma)$ for $p=1,...,q$. Then for almost every (in the sense of prevalence) Lipschitz map $f^{d}:\mathcal{Y}\rightarrow\mathbb{R}$ the delay observation map $f:=D_{q}[f^{d},\Phi^{\tau}]:\mathcal{Y}\rightarrow\mathbb{R}^{q}$ defined by
$y\mapsto\begin{bmatrix}f^{d}(y)&f^{d}(\Phi^{\tau}(y))&\ldots&f^{d}(\Phi^{(q-1)\tau}(y)\end{bmatrix}^{\top}$
is one-to-one on $\mathcal{A}$. The same holds for a set of $q$ distinct observables $f^{d}_{1},\ldots,f^{d}_{q}:\mathcal{Y}\to\mathbb{R}$, i.e.,
$f=\begin{bmatrix}f^{d}_{1}(y)&\ldots&f^{d}_{q}(y)\end{bmatrix}^{\top}.$
@@
The central message of the above theorem is that we can draw a close connection between the observable $f$ in the Koopman setting (Definition 1 and Remark 2) and the observation map $f^{d}$ in the embedding framework. The key statement for our purposes is that if the system dynamics of $\Phi^{\tau}$ is restricted to a compact set $\mathcal{A}$ with a finite dimension $d$, then we need to have at least $q>2d$ distinct measurements $f^{d}$—which jointly form the Koopman observable $f$—to obtain a one-to-one correspondence between $\Phi^{\tau}$ and $\varphi^{\tau}$:
$\Phi^{\tau}=E\circ\varphi^{\tau}\circ D_{q}[f^{d},\Phi^{\tau}]=E\circ\varphi^{\tau}\circ f,$
@@ 5 Combining equivariance and partial measurements
In fact, if we were able to obtain an exact finite-dimensional approximation of $\hat{\mathcal{K}}^{\tau}$, then we would quickly run into inconsistencies, as the same matrix would be a predictor for $(z_{1},z_{2},z_{3})$ and for $(z_{2},z_{3},z_{4})$. A quick calculation then shows that the prediction of any $z_{i}$ cannot depend on any other $z_{j}$, which would mean that $K^{\tau}$ is just a diagonal matrix. Our interpretation of this dilemma is that inconsistencies cannot be avoided, as in the generic situation, any finite-dimensional approximation is inexact, meaning that the local Koopman models have to be globally inconsistent or yield trivial dynamics. Note that this situation is different for other common observable functions such as Fourier modes. These have more sophisticated implications for the symmetry, as shown in Example 3.2
@@
As the above approach only works for control inputs that enter the original system in a purely linear fashion [NPP+23], an alternative approach is to accept that a purely linear approach may be too much to ask for. Instead, a coupling can be achieved by predicting the next state, project from $\Psi(z)$ onto the coordinates $z$, and then lift again for each system. Due to the project-then-lift step, we “synchronize” the local systems, at the cost of obtaining nonlinear dynamics. Similar observations in terms of trading the linearity for accuracy have been made in, e.g., [CALG24]. Note that if we use kernel EDMD with a radial basis function kernel, then this approach possesses a relation to Gaussian Process models.